# Consumption Partial Insurance

### v5

@@ 5.3.2 Results by age and position in the income distribution
Table 4 reports results from the quadratic specification (see table C.10 for the linear specification). $\phi^{(1)}$ decreases with age, so the average permanent shock affects consumption less as households age. In line with theory, this pattern reflects the increasing self-insurance role of financial wealth over the lifecycle. $\phi^{(2)}$, the transmission parameter of the typical shock away from the mean, also decreases with age. It drops from $0.006$ to $-0.008$ as we move from ages 30–40 to 41–50 (neither statistically significant) and becomes very negative and significant at $\phi^{(2)}=-0.104$ in the 51–65 bracket. As a result, large negative permanent shocks impact consumption more severely at older ages. Consider, again, $\zeta_{it}=-0.5$. At age 45, its pass-through is $\partial\Delta c_{it}/\partial\zeta_{it}=0.123$ (s.e.$=0.080$), while its pass-through at age 60 is almost double, $\partial\Delta c_{it}/\partial\zeta_{it}=0.218$ (s.e.$=0.077$).
@@
In our data, skewness of permanent shocks becomes more severe as households age, increasing in magnitude from $-0.65$ to $-2.39$ between ages 30–40 and 51–65 (table C.9). This is consistent with Guvenen et al. (2021), who document that skewness of earnings growth becomes more negative with the level of earnings and, consequently, with the earner’s age. There is more room for income to fall sharply in this case than among the young. The results in table 4, however, reveal that large bad shocks affect consumption more at older ages not only because those shocks become more extreme but also because the consumption response itself changes ($\phi^{(2)}$ becomes more negative). The type of events that underlie large income downturns at older ages, such as unemployment or declining health – see table C.3 for sources of bad shocks, also trigger permanent cuts of certain types of expenditure (food away from home, offspring education, or work-related expenses), which translate into a consumption response that differs between the old and the young. This does not mean that moderate bad shocks are worse insured among the old; for small or moderate shocks, the decreasing pattern in $\phi^{(1)}$ offsets the increasing magnitude of $|\phi^{(2)}|$.
@@
At face value, $\phi^{(2)}$ is more negative among the poor, and highly significant, suggesting that a given bad shock transmits more strongly among them. However, this provides only a partial picture of the underlying dynamics because the nature of income risk is vastly different between poor and rich. Appendix table C.9 shows that skewness of permanent shocks is positive among the poor, with a coefficient of $1.34$ (marginally insignificant), but negative among the rich, with a coefficient of $-4.86$ (multiple times larger than the baseline, and highly significant). One may thus study how large bad/good shocks from a given position in the underlying distributions, e.g., shocks in the 3rd or 97th percentiles, affect consumption in each case. Simulations are needed to infer the value of the specific shocks. [^44] We simulate a mixture of normal distributions separately for the poor and the rich, given the estimated income parameters of appendix table C.9. We provide details on similar simulations in appendix D.
Adopting this approach, the lower-tail shock shifts $c_{it}$ by $-0.017$ among the poor ($\Delta c_{it}=-0.017$, associated pass-through $0.129$) and by $-0.035$, twice as much, among the rich (pass-through $0.084$). By contrast, the upper-tail shock increases $c_{it}$ by $0.022$ among the poor (pass-through $0.075$) but by only $0.006$ among the rich (pass-through $0.057$). In other words, households in the top of the income distribution (the rich) exhibit a stronger consumption change in response to bad news, while households in the bottom (the poor) exhibit a stronger consumption change in response to good news.
@@
How does a less negative $\phi^{(2)}$ among the rich (smaller response to larger bad shocks) reconcile with a more negative $\phi^{(2)}$ among the old (stronger response), who are presumably also rich? The answer lies in the composition of the old (they are in fact evenly present in each income tertile) and in the interplay between income risk and consumption response. Skewness of permanent shocks is twice as large among the rich ($-4.86$) as among the old ($-2.39$). The consumption distribution is similar for both, so the rich must have overall more insurance to a given bad shock. This explains why $\phi^{(2)}$’s magnitude is four times smaller among them. In both cases (old, rich), bad permanent shocks from a given position in the distribution reduce consumption more than among their counterparts (young, poor).
@@ 5.3.3 Results by wealth and education
Notes: The table presents the estimates of the parameters of the quadratic consumption function, allowing them to vary across wealth and education groups. The underlying income moments also vary by wealth and education (results not shown for brevity). Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth over 19991999199919992019201920192019. No versus some college is defined on the basis of the highest level of education attained by the male spouse. All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses. Linear specification estimates appear in appendix table C.12.
@@ 5.3.4 Link with structural model and some estimates of welfare costs
Our quasi reduced-form approach so far has enabled us to estimate an asymmetric transmission of income shocks without relying on a particular parameterization of preferences, expectations, or income distribution. Estimating a fully specified model is outside the scope of this paper, but it is useful to evaluate our findings against the few structural lifecycle models that have recently incorporated higher-order income risk.
@@
We read these results as not too different from ours. Structural models seem to produce similar asymmetries in the transmission of income shocks, with some evidence that, as bad shocks hits, there is a limit to how much one can dissave or borrow. Therefore, such shocks end up affecting consumption more than other shocks. The magnitude of our pass-through rate of permanent shocks is uniformly lower than its structural counterpart (consumption transmission in the data likely reflects multiple insurance sources, beyond self-insurance and taxes examined in these models), yet other empirical studies using the updated PSID, e.g., ABB and Arellano et al. (2023), find similar rates to ours.
@@ Appendix A Derivation of the consumption function
This appendix shows how to derive the consumption function under a permanent-transitory specification for log income. We follow similar derivations in Blundell et al. (2008), BPP hereafter, though our approach is a refinement of theirs given the second-order approximation we carry out herein. We show below how to derive the most general (quadratic) consumption function (5); the linear function (4) is a special case of this. We let a time unit in the model correspond to one calendar year. However, data in the PSID in recent years are available only every second year, so we adapt the derivations to reflect this.
@@
[…]
@@ A.1 Taylor approximation to optimality conditions
The first term $\omega_{it+s+1}=\Gamma-\frac{1}{2}{\operatorname*{\mathbb{E}}}_{t}(\Delta\ln\lambda_{it+s+1}-\Gamma)^{2}$ captures the impact of interest rates, impatience, and precautionary motives on the growth of the marginal utility of wealth. To maintain tractability, we assume that $\omega$ is non-stochastic but possibly heterogeneous across households. The second term is an expectation error with ${\operatorname*{\mathbb{E}}}_{t}(\epsilon_{it+1})=0$; it captures idiosyncratic revisions to the marginal utility of wealth upon arrival of income shocks.
@@
where $\Delta c_{it+s}=\Delta\ln C_{it+s}-\theta_{it+s-1}\Delta(\mathbf{Z}_{it+s}^{\prime}\boldsymbol{\alpha})-\Xi_{it+s}$. We have split $\theta_{it+s-1}\omega_{it+s}$ into $\Xi_{it+s}$ and $\xi_{it+s}$; the first term reflects the gradient of consumption due to interest rates and impatience (discounting), the second term reflects unobserved consumption taste heterogeneity. We let demographics and time fixed effects pick up the effect of taste shifters and the gradient in the consumption path. Therefore $\Delta c_{it+s}$ can be obtained empirically from a regression of consumption growth on appropriate observables (demographics, time fixed effects).
@@ A.2 Taylor approximation to lifetime budget constraint
Let $G(\boldsymbol{\xi})=\ln\sum_{s=0}^{N}\exp\xi_{s}$ for $\boldsymbol{\xi}=(\xi_{0},\xi_{1},\dots,\xi_{N})^{\prime}$. Applying a second-order Taylor expansion of $G(\boldsymbol{\xi})$ around a deterministic $\boldsymbol{\xi}^{0}$, and taking expectations given information ${\cal I}$, yields
@@
$\xi_{s}$ =$=\begin{cases}\ln A_{it+s}&\text{ for }s=0\\ \ln Y_{it+s-1}-(s-1)\ln(1+r)\qquad\qquad&\text{ for }s=1,\dots,T-t+1\\ \end{cases}$
$\xi^{0}_{s}$ =$=\begin{cases}{\operatorname*{\mathbb{E}}}_{t-2}\ln A_{it+s}&\text{ for }s=0\\ {\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s-1}-(s-1)\ln(1+r)&\text{ for }s=1,\dots,T-t+1,\end{cases}$
@@
$+\frac{1}{2}\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}(1-\pi_{it})\vartheta_{it+s}^{Y}(\delta_{s\ell}-(1-\pi_{it})\vartheta_{it+\ell}^{Y}){\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})(\ln Y_{it+\ell}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+\ell})$
@@
The notation is: $\pi_{it}=\frac{Q_{1t}}{Q_{1t}+Q_{2t}}$ is equal to the expected share of financial wealth $Q_{1t}=\exp({\operatorname*{\mathbb{E}}}_{t-2}\ln A_{it})$ in the household’s total financial and human wealth at $t$, the latter defined as $Q_{2t}=\sum_{\kappa=0}^{T-t}\exp\left({\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+\kappa}-\kappa\ln(1+r)\right)$, i.e., the household’s expected lifetime income for the remaining of life; $\vartheta_{it+s}^{Y}=\exp\left(\operatorname*{\mathbb{E}}_{t-2}\ln Y_{it+s}-s\ln(1+r)\right)/Q_{2t}$ is an annuitization factor equal to the expected share of time $t+s$ household income in total lifetime income. $\pi_{it}$ and $\vartheta_{it+s}^{Y}$ pertain to expectations at $t-2$ so they are both known at $t-2$ and later.
@@
Using the income process and assuming the deterministic observables away (they cancel out in a first difference in expectations at a later stage), we reach $\ln Y_{it+s}=\ln Y_{it-2}+\sum_{\tau=-1}^{s}\zeta_{it+\tau}+v_{it+s}-v_{it-2}$, therefore $\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it-2}=\sum_{\tau=-1}^{s}\zeta_{it+\tau}+v_{it+s}$. It follows that $\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}{\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})=\zeta_{it}+\zeta_{it-1}+\vartheta_{it}^{Y}v_{it}$ because ${\operatorname*{\mathbb{E}}}_{t}\zeta_{it+\kappa}={\operatorname*{\mathbb{E}}}_{t}v_{it+\kappa}=0$ for $\kappa>0$, and $\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}=1$ by construction. The second-order term in income is more involved. This is given by
$\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}(1-\pi_{it})\vartheta_{it+s}^{Y}(\delta_{s\ell}-(1-\pi_{it})\vartheta_{it+\ell}^{Y}){\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})(\ln Y_{it+\ell}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+\ell})$
$=\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}(1-\pi_{it})\vartheta_{it+s}^{Y}(\delta_{s\ell}-(1-\pi_{it})\vartheta_{it+\ell}^{Y}){\operatorname*{\mathbb{E}}}_{t}\left((\zeta_{it}+\zeta_{it-1})^{2}+(\zeta_{it}+\zeta_{it-1})(v_{it+s}+v_{it+\ell})+v_{it+s}v_{it+\ell}\right)$
$=(1-\pi_{it})\pi_{it}(\zeta_{it}+\zeta_{it-1})^{2}+2\pi_{it}(1-\pi_{it})\vartheta_{it}^{Y}(\zeta_{it}+\zeta_{it-1})v_{it}+(1-\pi_{it})\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})v_{it}^{2}.$
@@
We derive this making use of $\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}=1$, $\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{Y}\vartheta_{it+\ell}^{Y}=(\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y})(\sum_{\ell=0}^{T-t}\vartheta_{it+\ell}^{Y})=1$, and the expectation of future shocks being zero by construction.
@@
Plugging these results in the approximation of the previous page, noting that assets at $t$ are ‘beginning-of-period’ and thus independent of the time $t$ income shocks, and taking a first difference in expectations between $t-2$ and $t$ yields

### v6

@@ 5.3.2 Results by age and position in the income distribution
Table 4 reports results from the quadratic specification (see table C.10 for the linear specification). $\phi^{(1)}$ decreases with age, so the average permanent shock affects consumption less as households age. In line with theory, this pattern reflects the increasing self-insurance role of financial wealth over the lifecycle. $\phi^{(2)}$, the transmission parameter of the typical shock away from the mean, also decreases with age. It drops from $0.006$ to 0.008-0.008 as we move from ages 30–40 to 41–50 (neither statistically significant) and becomes very negative and significant at $\phi^{(2)}=-0.104$ in the 51–65 bracket. As a result, large negative permanent shocks impact consumption more severely at older ages. Consider, again, $\zeta_{it}=-0.5$. At age 45, its pass-through is $\partial\Delta c_{it}/\partial\zeta_{it}=0.123$ (s.e.=0.080=0.080), while its pass-through at age 60 is almost double, $\partial\Delta c_{it}/\partial\zeta_{it}=0.218$ (s.e.=0.077=0.077).
@@
In our data, skewness of permanent shocks becomes more severe as households age, increasing in magnitude from 0.65-0.65 to 2.39-2.39 between ages 30–40 and 51–65 (table C.9). This is consistent with Guvenen et al. (2021), who document that skewness of earnings growth becomes more negative with the level of earnings and, consequently, with the earner’s age. There is more room for income to fall sharply in this case than among the young. The results in table 4, however, reveal that large bad shocks affect consumption more at older ages not only because those shocks become more extreme but also because the consumption response itself changes ($\phi^{(2)}$ becomes more negative). The type of events that underlie large income downturns at older ages, such as unemployment or declining health – see table C.3 for sources of bad shocks, also trigger permanent cuts of certain types of expenditure (food away from home, offspring education, or work-related expenses), which translate into a consumption response that differs between the old and the young. This does not mean that moderate bad shocks are worse insured among the old; for small or moderate shocks, the decreasing pattern in $\phi^{(1)}$ offsets the increasing magnitude of $|\phi^{(2)}|$.
@@
At face value, $\phi^{(2)}$ is more negative among the poor, and highly significant, suggesting that a given bad shock transmits more strongly among them. However, this provides only a partial picture of the underlying dynamics because the nature of income risk is vastly different between poor and rich. Appendix table C.9 shows that skewness of permanent shocks is positive among the poor, with a coefficient of $1.34$ (marginally insignificant), but negative among the rich, with a coefficient of 4.86-4.86 (multiple times larger than the baseline, and highly significant). One may thus study how large bad/good shocks from a given position in the underlying distributions, e.g., shocks in the 3rd or 97th percentiles, affect consumption in each case. Simulations are needed to infer the value of the specific shocks. [^44] We simulate a mixture of normal distributions separately for the poor and the rich, given the estimated income parameters of appendix table C.9. We provide details on similar simulations in appendix D.
Adopting this approach, the lower-tail shock shifts $c_{it}$ by 0.017-0.017 among the poor ($\Delta c_{it}=-0.017$, associated pass-through $0.129$) and by 0.035-0.035, twice as much, among the rich (pass-through $0.084$). By contrast, the upper-tail shock increases $c_{it}$ by $0.022$ among the poor (pass-through $0.075$) but by only $0.006$ among the rich (pass-through $0.057$). In other words, households in the top of the income distribution (the rich) exhibit a stronger consumption change in response to bad news, while households in the bottom (the poor) exhibit a stronger consumption change in response to good news.
@@
How does a less negative $\phi^{(2)}$ among the rich (smaller response to larger bad shocks) reconcile with a more negative $\phi^{(2)}$ among the old (stronger response), who are presumably also rich? The answer lies in the composition of the old (they are in fact evenly present in each income tertile) and in the interplay between income risk and consumption response. Skewness of permanent shocks is twice as large among the rich (4.86-4.86) as among the old (2.39-2.39). The consumption distribution is similar for both, so the rich must have overall more insurance to a given bad shock. This explains why $\phi^{(2)}$’s magnitude is four times smaller among them. In both cases (old, rich), bad permanent shocks from a given position in the distribution reduce consumption more than among their counterparts (young, poor).
@@ 5.3.3 Results by wealth and education
Notes: The table presents the estimates of the parameters of the quadratic consumption function, allowing them to vary across wealth and education groups. The underlying income moments also vary by wealth and education (results not shown for brevity). Low (high) wealth is defined on the basis of household wealth being less (more) than median real wealth over 1999199920192019. No versus some college is defined on the basis of the highest level of education attained by the male spouse. All columns use consumption data internally available in the PSID. Estimation is via equally weighted GMM; GMM standard errors clustered at the household level are in parentheses. Linear specification estimates appear in appendix table C.12.
@@ 5.3.4 Link with structural model and some estimates of welfare costs
Our quasi reduced-form approach so far has enabled us to estimate an asymmetric transmission of income shocks without relying on a particular parameterization of preferences, expectations, or the income distribution. Estimating a fully specified model is outside the scope of this paper, but it is useful to evaluate our findings against the few structural lifecycle models that have recently incorporated higher-order income risk.
@@
We read these results as not too different from ours. Structural models seem to produce similar asymmetries in the transmission of income shocks, with some evidence that, as bad shocks hits, there is a limit to how much one can dissave or borrow. The magnitude of our pass-through rate of permanent shocks is uniformly lower than its structural counterpart (consumption transmission in the data likely reflects multiple insurance sources, beyond self-insurance and taxes examined in these models), yet other empirical studies using the updated PSID, e.g., ABB and Arellano et al. (2023), find similar rates to ours.
@@
6 An alternative income process: persistent shocks
Our results so far rely on a permanent-transitory income process that leaves no room for shocks to have persistent, yet not fully permanent, effects on income. This unit root restriction in income may be perceived as unappealing in the context of biennial data and non-Gaussian shocks, especially if the various moments of shocks, on which the partial insurance estimates depend, are sensitive to it. Moreover, Kaplan and Violante (2010) show that BPP’s method of measuring insurance to permanent shocks is biased in the presence of borrowing limits, but the bias lessens if the framework is adapted to measure insurance to persistent shocks. This is unlikely to explain the low magnitude of $\phi^{(1)}$ that we find here (the bias tends to increase the linear pass-through rate), but it is unclear how it might affect $\phi^{(2)}$. For both these reasons, we relax the unit root restriction in income in this section.
Let persistent income $P_{it}$ follow an AR(1) process rather than a unit root. We may write this as $P_{it}=\rho P_{it-1}+\zeta_{it}$, where $0<\rho\leq 1$ measures income persistence. All other assumptions (including the distributions of income shocks) remain similar to the baseline. This modeling choice gives rise to the persistent-transitory formulation $\Delta y_{it}=(\rho-1)P_{it-1}+\zeta_{it}+\Delta v_{it}$ and the quasi-difference
$\Delta_{\rho}y_{it}=\zeta_{it}+\Delta_{\rho}v_{it},$ (9)
where $\Delta_{\rho}x_{t}=x_{t}-\rho x_{t-1}$ and $\zeta$ reflects a persistent rather than a permanent shock. Clearly, the permanent-transitory process is a special case of this for $\rho=1$.
Despite the modification in the income process, the specification of the linear and quadratic consumption functions remains unchanged from the baseline processes (4) and (5), respectively. However, the structure that underlies the transmission parameters $\phi^{(1)}$, $\phi^{(2)}$, etc, explicitly involves $\rho$ in this case. Appendix A details the derivation of the consumption function with income persistence and some further adjustments when the data are observed biennially.
The identification strategy for the income process must now be adjusted to account for the AR(1) component and the presence of an additional parameter, $\rho$. Formally, the variance of shocks and $\rho$ are jointly identified through
$\sigma_{\zeta_{t}}^{2}$ $={\operatorname*{\mathbb{E}}}(\Delta_{\rho}y_{it}\times(\rho\Delta_{\rho}y_{it-1}+\Delta_{\rho}y_{it}+\rho^{-1}\Delta_{\rho}y_{it+1})),$
$\sigma_{v_{t}}^{2}$ $=-\rho^{-1}{\operatorname*{\mathbb{E}}}(\Delta_{\rho}y_{it}\times\Delta_{\rho}y_{it+1}),$
${\operatorname*{\mathbb{E}}}(\Delta_{\rho}y_{it}\times\Delta y_{it-2})$ $={\operatorname*{\mathbb{E}}}((y_{it}-\rho y_{it-1})\times(y_{it-2}-y_{it-3}))=0.$
The first two moments appear slightly modified also in the baseline; the third moment is new and pins down $\rho$. An additional time period of data is needed (t3t-3, for a minimum of 5 periods in total) to separately identify income persistence from the variance of persistent shocks. The skewness and kurtosis of shocks are then identified from similarly modified higher-order income moments that mimic the baseline. Identification of the consumption transmission parameters is analogous: the quasi-difference $\Delta_{\rho}y_{it}$ replaces $\Delta y_{it}$, and $\sum_{\tau=-1}^{1}\rho^{-\tau}\Delta_{\rho}y_{it+\tau}$ replaces the long sum $\sum_{\tau=-1}^{1}\Delta y_{it+\tau}$ in all previous identifying statements, but otherwise the strategy is similar to the baseline. Appendix B shows the identifying statements in detail, including the adjustments needed for biennial data and measurement error; identification in the quadratic case again obtains for any distribution of income shocks except Bernoulli.
Table 7: Estimates of the consumption function, income persistence
Consumption fn.: Quadratic
Income moments: 2nd, 3rd, 4th
Income persistence: $\rho=0.931$ (s.e.=0.028=0.028)
Consumption data: PSID
$\phi^{(1)}$ 0.112 (0.024)
$\psi^{(1)}$ -0.010 (0.066)
$\phi^{(2)}$ -0.049 (0.026)
$\psi^{(2)}$ 0.019 (0.037)
$\omega^{(22)}$ 0.430 (0.847)
$\sigma^{2}_{\xi}$ 0.019 (0.001)
$\sigma^{2}_{u_{c}}$ 0.044 (0.002)
Notes: The table presents the estimates of the parameters of the quadratic consumption function, when income features linear persistence. Estimation is via equally weighted GMM; block bootstrap standard errors are in parentheses. The estimates of the income process appear in appendix table C.13; linear consumption specification estimates appear in appendix table C.14.
Appendix table C.13 reports the estimates of the income process from this specification. Three points emerge. First, when targeting only the second moments of income, we estimate $\rho=0.846$ (s.e.=0.038=0.038), which lies at the lower end of persistence estimates in the literature but within ABB’s nonlinear persistence range of 0.5–1 over a similar period. The biennial nature of the data likely influences this, even though $\rho$ reflects annual persistence. As $\rho$ is below its baseline value, the variance of persistent shocks is higher than the baseline to help match the variance of income growth in the data (yet the transitory variance is smaller). Second, adding third-order moments yields negative skewness for both shocks but leaves the estimates of $\rho$ and the variances unchanged. Third, targeting all moments up to fourth-order, we estimate $\rho=0.931$ (s.e.=0.028=0.028). The estimates of the variances align with the baseline values of table 2, as also do the skewness coefficients at -0.91 and -1.178 for the two types of shocks, respectively. We estimate kurtosis at 28.891 and 52.091; for persistent shocks, this is lower than the baseline, yet both distributions remain highly leptokurtic and left-skewed despite the AR(1) component.
Table 7 reports the estimates of the quadratic consumption function (see table C.14 for the linear specification). We estimate the transmission parameter of the average persistent shock at $\phi^{(1)}=0.112$, somewhat lower than the value 0.134 when shocks are permanent. The average persistent shock is therefore slightly better insured than the average permanent shock. Kaplan and Violante (2010) similarly observe a drop in the pass-through rate (an increase in partial insurance) when $\rho$ falls and shocks become less permanent. We further estimate the transmission parameter of persistent shocks away from the middle of the distribution at $\phi^{(2)}=-0.049$, negative and larger in magnitude than the baseline value of -0.04. Negative persistent shocks remain more unsettling than positive ones: they are associated with a larger pass-through rate into consumption, one that increases with the severity of the shock. The asymmetric response to bad versus good income news is therefore qualitatively and numerically robust to the earlier unit root assumption. Finally, transitory shocks have economically small and statistically zero pass-through rates. [^47] In the context of the linear consumption specification (appendix table table C.14), we estimate $\phi^{(1)}=0.089$ when we target only second moments, which gradually increases to match its quadratic specification counterpart as we introduce higher-order moments. This is driven by the increase in the value of $\rho$ as we add higher moments. Kaplan and Violante (2010) also find lower pass-through rates (more insurance) at lower values of $\rho$. Transitory shocks remain fully insured, statistically., [^48] It is straightforward to further condition the results on age, position in the income distribution, wealth and education. We do not report these results for brevity.
@@ Appendix A Derivation of the consumption function
This appendix shows how to derive the consumption function under the general AR(1) specification for log income in section 6; the consumption function based on the baseline permanent-transitory income process is a special case of this. We follow similar derivations in Blundell et al. (2008), BPP hereafter, though our approach is a refinement of theirs given the second-order approximation we carry out herein and the relaxation of the unit root assumption for income. We show below how to derive the most general (quadratic) consumption function; the linear function is a special case of this. We let a time unit in the model correspond to one calendar year. However, data in the PSID in recent years are available only every second year, so we adapt the derivations to reflect this.
@@
$[C_{it+s}]$ $:{}$ $\quad\widetilde{U}_{C}(\widetilde{C}_{it+s})\exp(-\mathbf{Z}_{it+s}^{\prime}\boldsymbol{\alpha})$ $=\lambda_{it+s}$
$[A_{it+s+1}]$ $:{}$ $\quad\beta(1+r){\operatorname*{\mathbb{E}}}_{t+s}\lambda_{it+s+1}$ $=\lambda_{it+s},$
@@ A.1 Taylor approximation to optimality conditions
The first term $\omega_{it+s+1}=\Gamma-\frac{1}{2}{\operatorname*{\mathbb{E}}}_{t}(\Delta\ln\lambda_{it+s+1}-\Gamma)^{2}$ captures the impact of the interest rate, impatience, and precautionary motives on the growth of the marginal utility of wealth. To maintain tractability, we assume that $\omega$ is non-stochastic but possibly heterogeneous across households. The second term is an expectation error with ${\operatorname*{\mathbb{E}}}_{t}(\epsilon_{it+1})=0$; it captures idiosyncratic revisions to the marginal utility of wealth upon arrival of income shocks.
@@
where $\Delta c_{it+s}=\Delta\ln C_{it+s}-\theta_{it+s-1}\Delta(\mathbf{Z}_{it+s}^{\prime}\boldsymbol{\alpha})-\Xi_{it+s}$. We have split $\theta_{it+s-1}\omega_{it+s}$ into $\Xi_{it+s}$ and $\xi_{it+s}$; the first term reflects the gradient of consumption due to impatience (discounting) and the interest rate, the second term reflects unobserved consumption taste heterogeneity. We let demographics and time fixed effects pick up the effect of taste shifters and the gradient in the consumption path. Therefore $\Delta c_{it+s}$ can be obtained empirically from a regression of consumption growth on appropriate observables (demographics, time fixed effects).
@@ A.2 Taylor approximation to lifetime budget constraint
[…]
@@
[…]
$\xi^{0}_{s}$ $=\begin{cases}{\operatorname*{\mathbb{E}}}_{t-2}\ln A_{it+s}&\text{ for }s=0\\ {\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s-1}-(s-1)\ln(1+r)&\text{ for }s=1,\dots,T-t+1,\end{cases}$
@@
$+\frac{1}{2}(1-\pi_{it})\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{Y}(\delta_{s\ell}-(1-\pi_{it})\vartheta_{it+\ell}^{Y}){\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})(\ln Y_{it+\ell}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+\ell})$
@@
The notation is: $\pi_{it}=\frac{Q_{1t}}{Q_{1t}+Q_{2t}}$ is equal to the expected share of financial wealth $Q_{1t}=\exp({\operatorname*{\mathbb{E}}}_{t-2}\ln A_{it})$ in the household’s total financial and human wealth at $t$, the latter defined as $Q_{2t}=\sum_{\kappa=0}^{T-t}\exp\left({\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+\kappa}-\kappa\ln(1+r)\right)$, i.e., the household’s expected lifetime income for the remaining of life; $\vartheta_{it+s}^{Y}=\exp\left(\operatorname*{\mathbb{E}}_{t-2}\ln Y_{it+s}-s\ln(1+r)\right)/Q_{2t}$ is an annuitization factor equal to the expected share of time $t+s$ household income in total lifetime income. $\pi_{it}$ and $\vartheta_{it+s}^{Y}$ pertain to expectations at t2t-2 so they are both known at t2t-2 and later.
@@
Using the income process and assuming the deterministic observables away (they cancel out in a first difference in expectations at a later stage), we reach $\ln Y_{it+s}=\rho^{s+2}\ln Y_{it-2}+\sum_{\tau=-1}^{s}\rho^{s-\tau}\zeta_{it+\tau}+v_{it+s}-\rho^{s+2}v_{it-2}$, therefore $\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s}=\sum_{\tau=-1}^{s}\rho^{s-\tau}\zeta_{it+\tau}+v_{it+s}$, given that $s\geq-1$. It follows that $\sum_{s=0}^{T-t}\vartheta_{it+s}^{Y}{\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})=\Theta_{it}^{Y}(\zeta_{it}+\rho\zeta_{it-1})+\vartheta_{it}^{Y}v_{it}$, where $\Theta_{it}^{Y}=\sum_{\kappa=0}^{T-t}\vartheta_{it+\kappa}^{Y}\rho^{\kappa}$, because ${\operatorname*{\mathbb{E}}}_{t}\zeta_{it+\kappa}={\operatorname*{\mathbb{E}}}_{t}v_{it+\kappa}=0$ for $\kappa>0$. The second-order term in income is more involved. This is given by
$\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{Y}(\delta_{s\ell}-(1-\pi_{it})\vartheta_{it+\ell}^{Y}){\operatorname*{\mathbb{E}}}_{t}(\ln Y_{it+s}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+s})(\ln Y_{it+\ell}-{\operatorname*{\mathbb{E}}}_{t-2}\ln Y_{it+\ell})$
$=\sum_{s=0}^{T-t}\sum_{\ell=0}^{T-t}\vartheta_{it+s}^{Y}(\delta_{s\ell}-(1-\pi_{it})\vartheta_{it+\ell}^{Y}){\operatorname*{\mathbb{E}}}_{t}\left(\rho^{s}\rho^{\ell}(\zeta_{it}+\rho\zeta_{it-1})^{2}+(\zeta_{it}+\rho\zeta_{it-1})(\rho^{s}v_{it+\ell}+\rho^{\ell}v_{it+s})+v_{it+s}v_{it+\ell}\right)$
$=\widetilde{\pi}_{it}(\zeta_{it}+\rho\zeta_{it-1})^{2}+2\widetilde{\widetilde{\pi}}_{it}\vartheta_{it}^{Y}(\zeta_{it}+\rho\zeta_{it-1})v_{it}+\vartheta_{it}^{Y}(1-(1-\pi_{it})\vartheta_{it}^{Y})v_{it}^{2},$
@@
where $\widetilde{\pi}_{it}=\widetilde{\Theta}_{it}^{Y}-(1-\pi_{it})(\Theta_{it}^{Y})^{2}$, $\widetilde{\Theta}_{it}^{Y}=\sum_{\kappa=0}^{T-t}\vartheta_{it+\kappa}^{Y}\rho^{2\kappa}$, and $\widetilde{\widetilde{\pi}}_{it}=1-(1-\pi_{it})\Theta_{it}^{Y}$. We derive this noting that the expectation of future shocks is zero by construction.
@@
Plugging these results in the approximation of the previous page, noting that assets at $t$ are ‘beginning-of-period’ and thus independent of the time $t$ income shocks, and taking a first difference in expectations between t2t-2 and $t$ yields